OpenStudy (anonymous):

modulus of 2(cos(pi/3), isin(pi/3)

OpenStudy (anonymous):

Help please!

OpenStudy (anonymous):

Are you asking what the remainder is of 2(cos(pi/3) divided by isin(pi/3) ?

OpenStudy (anonymous):

no I am asking of the MODULUS of this equation.

OpenStudy (anonymous):

Ahhh. Ok. So then modulus = sqrt( (2(cos(pi/3))^2 + (sin(pi/3))^2 ) right?

OpenStudy (anonymous):

no this modulus is of complex numbers

OpenStudy (anonymous):

modulus and argument of complex numbers..get it?

OpenStudy (anonymous):

Right. Let me make sure I understand what you are asking. The number is 2cos(pi/3) + isin(pi/3) ? Or did you give two separate numbers. As far as I understand, the modulus of a complex numebr z = a + bi is |z| = sqrt( a^2 + b^2)